INAF queue observing, 2017 Sep 28-29

Observer: A. Rossi, R. Carini
Telescope Operator: G. Bechetti
Support Astronomer: O. Kuhn
Instrument: LUCI1, LUCI2

Summary

At the sunset the sky is almost clear but the humidity is about 85%.

Night duration (12\xB0 el twilight) 2:00am- 12:20pm UT = 10 h 20m

Observing time: 1 hrs 20 m (13%)

Weather time loss: 9 h (87%)

technical time loss * *h (0%)

Data Summary

Project_RUN Objects Instrument Filter ExpTime Notes
Antoniucci8 V733 telluric LUCI zJ+HK   ok
Antoniucci8 V733 LUCI zJ+HK 80x4 in zJ, 30x4 in HK completed
Antoniucci8 V733 LUCI J 25s completed

Details

1:45 begin the preset for the observation of the Telluric star for V733, a Antoniucci program *02:07 collimation terminated, start acq
  • acq luci1.20170929.0001-0003
  • acq luci2.20170929.0005-0007
  • sci luci1.20170929.0004-0011
  • sci luci2.20170929.0008-0015
  • seeing about 0.8''

02:29 slew to the scientific target V733
  • luci2 doesn't find the guide star
  • 02:34 load again the script
  • 02:37 problem script go stuck on LUCI 1 before slit acq image
  • 02:42 start again, 5 min lost
  • acq luci1.20170929.0012-0015
  • acq luci2.20170929.0016-0019
  • seeing 0.9 on dimm
  • sci luci1.20170929.0016-0019 zJ
  • sci luci2.20170929.0020-0023 zJ
  • seeing ~1 on dimm at 03:00
  • sci luci1.20170929.0020-0023 HK
  • sci luci2.20170929.0024-0027 HK

03:04 slew to imaging ob in J for V733
  • seeing ~1 on guider at 3:05
  • sci luci1.20170929.0024-0028
  • sci luci2.20170929.0028-0032

03:13 slew to PVCep
  • preset error at 3:14
  • clouds at 3:18 we stop

04:00 the humidity rises to 95% and the temperature to the dew point. TO close the dome

calibrations Antoniucci** flat zj luci1.0029-0034, luci2.0033-0037 there are problems with the configurations. With LUCI2 the absorption lines are at the center instead at the right. We have taken an arc with the Argon lamp, and the lines in LUCI2 are shifted respect the correct position. The same happens with the Neon lamp

* problem solved* at 12:30: o need to move on and off the mirror
  • HKspec Antoniucci 0.75" slit flats USELESS BECAUSE LIGHTS WERE ON TO WORK ON LBC
  • luci1.20170929.0060-0064
  • luci2.20170929.0080-0084
  • HKspec Antoniucci 0.75"arcs ARGON OK
  • luci1.20170929.0065-0069
  • luci2.20170929.0085-0089
  • HKspec Antoniucci 0.75"arcs NEON OK
  • luci1.20170929.0070-0074
  • luci2.20170929.0090-0094
  • HKspec Antoniucci 0.75"arcs XENON OK
  • luci1.20170929.0075-0076
  • luci2.20170929.0095-0096
  • HKspec Antoniucci 0.75" slit flats AGAIN OK FINALLY DONE
  • luci1.20170929.0077-0081
  • luci2.20170929.0097-0101
  • J flats Antoniucci imaging DONE
  • luci1.20170929.0082-0086
  • luci2.20170929.0102-0106

-- %USERSIG{RobertaCarini - 2017-09-29}%

Comments

 
Topic revision: r9 - 30 Sep 2017, RobertaCarini
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback